# Model-Based Soft Maximization of Suitable Metrics of Long-Term Human Power

### v1

@@ 2.2 Bounded Trade-Off Aggregation
How should $r$ aggregate all humans’ ICCEA power $W_{h}(s)$ to determine its own intrinsic reward $U_{r}(s)$? This depends on what incentives we want to give $r$ regarding changes in (i) the inter-human power distribution, (ii) the inter-temporal power distribution, and (iii) the power distribution across different realizations of uncertainty. Similar questions abound in welfare theory, from which we benefit here. Requiring anonymity, continuity, strict monotonicity in each $W_{h}(s)$, and an independence axiom (“Independence of unconcerned agents”) implies we must again use a separable form, $U_{r}(s)=F^{H}(\sum_{h}f^{H}(W_{h}(s)))$, with continuous, strictly increasing $f^{H}$ and $F^{H}$ (Fleming 1952).
[…]
Concerning (ii), we find it natural to also disincentivize $r$ from trading off $\mathcal{H}$’s current power against $\mathcal{H}$’s later power too much: $r$ should generally prefer a trajectory $(s_{1},s_{2})$ where $W_{h}(s_{1})=W_{h^{\prime}}(s_{2})=1$ and $W_{h}(s_{2})=W_{h^{\prime}}(s_{1})=2$ to a trajectory $(s_{1},s_{2})$ where $W_{h}(s_{1})=W_{h^{\prime}}(s_{1})=1$ and $W_{h}(s_{2})=W_{h^{\prime}}(s_{2})=2$. This means that $F^{H}$ should be strictly concave, which incentivizes a rather homogenous power distribution along time. Since $f^{H}(w)<0$, $F^{H}$ needs to be defined for negative values only. Finally, we motivate our concrete choice of $F^{H}$ and $\pi_{r}$ (and of $f^{G}$) by the following independence requirement. Assume we introduce an additional uncertainty into the world model (e.g., a formerly not modelled change of overall circumstances) whose consequence is that all goal attainment probabilities $V^{e}_{h}(s,g_{h})$ are multiplied by some common factor $b\in(0,1)$. Then $\pi_{r}$ shall not change. The simplest way to fulfil this requirement is to put $f^{G}(v)=v^{\zeta}$ (see above), $F^{H}(y)=-(-y)^{\eta}$ with $\eta>1$, [^5] An $\eta\approx 1$ makes approximate distributed computing of $V_{r}$ for rarely interacting subpopulations easier (see Supplement). and to use either an argmax policy for $\pi_{r}$ or a power-law-like soft policy with $\pi_{r}(s)(a)\propto(-Q_{r}(s,a_{r}))^{-\beta_{r}}$ for some $\beta_{r}>0$ (the minus signs are needed since $Q_{r}(s,a_{r})<0$). Choosing $\beta_{r}<\infty$ will allow the robot some exploration, e.g., to improve its world model, and will help avoiding remaining safety risks from misalignment with “true” power maximization, i.e., driving relevant aspects of ‘power’ that our metrics do not cover into very undesirable states, similar to (Zhuang and Hadfield-Menell 2020).
Re (iii), standard axioms imply taking expectations in eqns. (6) and (11). To increase $r$’s corrigibility, we assume goals might change anytime and take i.i.d. $g(t)\sim\mathcal{G}$ in (6).
Table 1 summarizes all design choices so far. The final set of equations for the fully observed case (see the Supplement for the partially observed case) is this:
[…]
$\pi_{r}(s)(a)$ $\propto(-Q_{r}(s,a_{r}))^{-\beta_{r}},$ (7)
For easier interpretation during analyses, one might eventually transform the quantities $X_{h}$, $U_{r}$ and $V_{r}$ back into units of bits: $W_{h}(s)=\log_{2}X_{h}(s)$ (individual power in $S$), $\operatorname{apow}(s)=-(\xi\eta)^{-1}\log_{2}(-U_{r}(s))$ (aggregate power in $s$), $\operatorname{tpow}(s)=-(\xi\eta)^{-1}\log_{2}(-V_{r}(s))$ (total power from $s$ on).
@@
In an acyclic environment, one can use backward induction to solve for the unique solution of (3)–(11) (see below). Otherwise, eqns. (3)–(5) define a continuous self-map $F$ on a finite-dimensional closed convex polytope of values $(Q^{m}_{h},\pi_{h},V^{m}_{h})$ and must thus have a solution due to Brouwer’s fixed point theorem. Due to the softmax, $F$ is neither a contraction nor is the resulting value iteration map monotonic unless $\beta_{h}<\beta_{h}^{1}$ for some $\beta_{h}^{1}>0$, so the solution can be non-unique in cyclic environments. Eqns. (6)–(11) also define a continuous self-map $G$ from a finite-dimensional bounded convex set $D$ of values $(Q_{r},\pi_{r},V^{e}_{h},X_{h},U_{r},V_{r})$ which is however not closed because $G$ is undefined for zero values in $Q_{r}$ or $X_{h}$. [^6] We could fix this by adding a small constant $\epsilon>0$ to $Q_{r}$ and $X_{h}$ before taking powers, which makes $G$ defined on a closed convex polytope as well, so that it must have a fixed point. Still, for $\beta_{r}=0$, there is a unique solution because $V^{e}_{h}$ is then the value function of a fixed policy, $U_{r}$ is independent of $V_{r}$, and so $V_{r}$ is also the value function of a fixed policy and reward function. Since $F,G$ are continuous in $\beta_{h},\beta_{r}$, we thus conjecture that homotopy / continuation methods can single out a unique “principal” solution for any $\beta_{h},\beta_{r}>0$ as in Goeree, Holt, and Palfrey (2016) even in cyclic environments.
@@ 3.1 Complex Multi-Agent Environments:
If $\mathcal{S}$ and $\mathcal{A}$ are large but $\mathcal{H}$ and $\mathcal{G}_{h}$ are not, $r$ can use a learning approach like the following. [^7] If also $\mathcal{H}$ and $\mathcal{G}_{h}$ are large, $r$ could use single neural networks $Q^{m}_{\mathcal{H}},\pi_{\mathcal{H}},V^{m}_{\mathcal{H}},V^{e}_{\mathcal{H}},X_{\mathcal{H}}$ that take feature vectors describing $h$ and $g_{h}$ as additional inputs, and use samples $h,g_{h}$ to train these networks and to approximate the sums in eqns. (9) and (10).
@@
For each $h\in\mathcal{H},g_{h}\in\mathcal{G}_{h}$ separately, learn tables or neural network approximations of $Q^{m}_{h}$ and $\pi_{h}$. Generate samples $(s,g_{h},a,s^{\prime})$ using a slowly updated $\beta_{h}$-softmax policy $\pi_{h}(s,g_{h})$ based on $Q_{h}^{m}$ with a decreasing amount of additional exploration, the prior policy $a_{-h}\sim\mu_{-h}(s)$ for other humans, and, to eventually learn the minimum in (3), an $\epsilon$-greedy policy for $a_{r}$ based on the negative expected value $-\operatorname*{\mathbb{E}}_{s^{\prime}\sim s,a}(U_{h}(s^{\prime},g_{h})+\gamma_{h}V^{m}_{h}(s^{\prime},g_{h}))$ with $\epsilon\to 0$. Then use expected SARSA targets on a time-scale faster than $\pi_{h}$.
@@
Based on the learned $\pi_{h}$, now aim to simultaneously learn tables or network approximations of $V^{e}_{h}$ and $X_{h}$ for all $h$, and of either $Q_{r}$ (DQN approach) or $V_{r}$ (actor-critic (AC) approach). Generate data samples $(s,g,a,s^{\prime})$ from rollouts using the fixed policies $\pi_{h}(s,g_{h})$, and either a $\beta_{r}^{\prime}$-softmax policy based on $Q_{r}$ with $\beta_{r}^{\prime}\nearrow\beta_{r}$, or a network approximation of $\pi_{r}$ trained on $V_{r}$ with entropy regularisation. Sample a new $g$ every $N_{g}$ steps. Instead of the direct expectation calculations of (6), (8), (9), then update the tables or networks via batched SGD using the update targets
@@ 4.1 Analysis of paradigmatic situations
We analyze the behavioral incentives of our objective in several paradigmatic situations in the Supplement.
@@
If the robot can make binding commitments (e.g. by offering labelled buttons causing different behaviors), make secret plans to react to $h$’s actions in certain ways (e.g. by offering unlabelled buttons), or act without waiting for button presses, it will generally make such commitments if it assumes $h$ to be sufficiently rational to use that information to act in ways that make $r$ do what $h$ wants. This is because $\pi_{h}$ does not depend on $\pi_{r}$ (encoding what $r$ plans) but on $\mathcal{A}_{r}$ (encoding what $r$ has committed to). Via such commitments, $r$ will establishing semantic connections between human actions (e.g., speech acts) and certain potentially complex behaviors of its own, to be able to act as an instruction-following assistant, somewhat similar to Reddy, Levine, and Dragan (2022).
@@
While an ‘empowerment’ maximizing robot would present humans with as many action options as possible, our robot will avoid overwhelming humans with too many options. If $r$ can choose to give $h$ any number $k$ of actions, each of which fulfils a different goal, it will choose $k\approx(e^{\beta_{h}}-1)/(\zeta-1)$ (noting that $\zeta>1$), since for larger $k$, $h$’s goal attainment probability will decrease due to bounded rationality so much that $h$’s effective power decreases despite the theoretical potential to fulfil more goals.
@@
The robot will sometimes ask back before obeying a command to perform an action because doing so decreases $h$’s error rate related to its bounded rationality. Also, if the action is irreversible, taking the action will remove $h$’s subsequent ability to make that choice, so delaying action leaves $h$ with decision power for longer. However, $r$ will eventually obey because otherwise $h$ would not have that choice in the first place. As can be expected, the number of times $r$ asks back increases with larger $\gamma_{h}$ and $\gamma_{r}$, i.e., the more patient $h$ and $r$ are, and decreases with larger $\beta_{h}$. E.g., in a minimal model, with $\gamma_{h}=0.99$, $\gamma_{r}=0$, and an effective human error rate of 10%, $r$ will already ask back twice before obeying a command.
@@
The robot will tend to follow human social norms that generally foster goal achievement. This is because $r$ will model most $h$ as expecting most others to follow the norm ($\mu_{-h}$), will thus expect those $h$ to also follow the norm ($\pi_{h}$) since that increases most $V^{m}_{h}(s,g_{h})$ for most goals. Thus $r$ will also follow the norm to prevent reducing those $h$’s power from harm or mis-coordination. If $r$ assumes humans to have internalized the norm into habits ($\pi^{0}_{h}$), it will even commit to following the norm itself to coordinate better in bilateral encounters without a social context.
@@
If the robot can split a total amount $M$ of resources between $h_{1}$ and $h_{2}$, and for $h_{i}$ to have resources $m$ translates into having a power of $W_{h}=f(m)$ bits, then $r$ will generally prefer an equal split, at least if $f$ is linear [^8] E.g., a linear $f$ seems plausible if $M$ is money that can be spent for paying others to make independent choices in one’s favor. , concave, or not too convex, and will only prefer an unequal split or even full resource concentration with substantially convex $f$. Larger $\xi$ will make the split more equal.
@@
The robot might choose to make humans have incorrect beliefs $\mu_{-h}$ about each others’ behavior in situations where $r$ can do so in the first place and where correct beliefs would lower effective goal attainments $V^{e}_{h}$, e.g. because most strategic (quantal response) equilibria are bad and most “social optima” (in terms of total power $U_{r}$) are far from strategic equilibrium.
@@
If $r$ cares for $h$’s future power ($\gamma_{r}\gg 0$) and can provide $h$ with the means to harm themselves, $r$ will trade off the temporary power increase from having these means against the possible later disempowerment from harm. So $r$ will provide the means only if it believes that $h$ is sufficiently rational and that $h$ will most probably not actually harm themselves.
@@
If $r$ has a pause and a destroy button and might disable either, it will generally enable only the pause button to prevent partially disempowering $h$ permanently by not being able to assist $h$ when destroyed. Only if $r$ thinks it very unlikely that $h$ will use the destroy button will it enable it to give $h$ this additional power. If $r$ thinks it is very helpful to $h$, it might even disable the pause button.

### v2

@@ 2.2 Bounded Trade-Off Aggregation
How should the robot aggregate all humans’ ICCEA power $W_{h}(s)$ to determine its own intrinsic reward $U_{r}(s)$? This depends on what incentives we want to give $r$ regarding changes in (i) the inter-human power distribution, (ii) the inter-temporal power distribution, and (iii) the power distribution across different realizations of uncertainty. Similar questions abound in welfare theory, guiding the way.
We again want anonymity, continuity, strict monotonicity in each $W_{h}(s)$, and an independence axiom (“Independence of unconcerned agents”). Again, this implies a separable form, $U_{r}(s)=F^{H}(\sum_{h}f^{H}(W_{h}(s)))$ with continuous, strictly increasing functions $f^{H}$ and $F^{H}$ (Fleming 1952).
Inter-human trade-offs
We do not want $r$ to concentrate power in the hands of a few, so we require the Pigou–Dalton principle of inequality aversion (Pigou 1912). This implies $f^{H}$ must be strictly concave. The functions most commonly used in such a context are those of “constant relative” and “constant absolute inequality aversion” (Amiel, Creedy, and Hurn 1999). As $W_{h}(s)=\log_{2}X_{h}(s)$, a natural choice is to use constant absolute inequality aversion w.r.t. $W_{h}(s)$ since that is equivalent to constant relative inequality aversion w.r.t. $X_{h}(s)$. This implies that $f^{H}(W_{h}(s))=-2^{-\xi W_{h}(s)}=-X_{h}(s)^{-\xi}$ for some $\xi>0$.
As it turns out, if we put $\xi=1$, we disincentivize $r$ from “taking away a person’s last binary choice” in the sense that reducing one human’s $W_{h}(s)$ from one bit to zero bits cannot be made up by increasing any other human’s $W_{h^{\prime}}(s)$ from $\geq 1$ bit to any value, because $2^{-1}-2^{-W_{h^{\prime}}(s)}\geq-1>-2^{-0}-2^{-w}$ for any $w\geq 1$. This is related closely to the idea of minimal individual rights (Pattanaik and Suzumura 1996). We thus choose $\xi=1$. [^4] To get a feeling for the effects of this choice of $f^{H}$ in a deterministic environment: for large enough $k$, taking away one of $k$ options from $h$ can be compensated by either giving at least two additional options to another $h^{\prime}$ with $k$ options or at least five additional options to some $h^{\prime}$ with $2k$ options.
Intertemporal trade-offs
We find it natural to also disincentivize $r$ from trading off current vs. later human power too much. It should generally prefer a trajectory $(s_{1},s_{2},\dots)$ with a rather homogenous power distribution along time, such as $W_{h}(s_{1})=W_{h^{\prime}}(s_{2})=1$ and $W_{h}(s_{2})=W_{h^{\prime}}(s_{1})=2$, to a trajectory where, say, $W_{h}(s_{1})=W_{h^{\prime}}(s_{1})=1$ and $W_{h}(s_{2})=W_{h^{\prime}}(s_{2})=2$. This means that $F^{H}$ should be strictly concave. Note that since $f^{H}(w)<0$, $F^{H}$ needs to be defined for negative values only.
We motivate our concrete choice of $F^{H}$ and $\pi_{r}$ (and of $f^{G}$) by the following independence requirement. Assume we introduce an additional uncertainty into the world model (e.g., a formerly not modelled change of overall circumstances) whose consequence is that all goal attainment probabilities $V^{e}_{h}(s,g_{h})$ are multiplied by some common factor $b\in(0,1)$. Then this should not change the policy $\pi_{r}$. The simplest way to fulfil this is to put $f^{G}(v)=v^{\zeta}$ (as done already above), $F^{H}(y)=-(-y)^{\eta}$ with $\eta>1$, and to use either an argmax policy for $\pi_{r}$ or a power-law-like policy with $\pi_{r}(s)(a)\propto(-Q_{r}(s,a_{r}))^{-\beta_{r}}$ for some $\beta_{r}>0$. Note that the minus signs are needed since $Q_{r}(s,a_{r})<0$.
We choose $\beta_{r}<\infty$ to allow the robot some exploration, e.g., to improve its world model. This should also help avoiding remaining safety risks when our metric misses some subtle but important aspects of ‘power’ that might thus be driven to very undesirable states under a full maximization of $V_{r}$, similar to (Zhuang and Hadfield-Menell 2020).
Aggregation across uncertainty
To deal with uncertain successor states $s^{\prime}$, standard axioms suggest we should simply take expectations in eqns. (4) and (9). When dealing with uncertain goals $g_{h}$, we want $r$ to be rather corrigible. So we assume goals might change anytime and take an expectation over independent uniform draws $g_{h}$.
Table 2 summarizes all the above design choices.
@@
In an acyclic environment, one can use backward induction to solve for the unique solution of (1)–(9) (see below). Otherwise, eqns. (1)–(3) define a continuous self-map $F$ on a finite-dimensional closed convex polytope of values $(Q^{m}_{h},\pi_{h},V^{m}_{h})$ and must thus have a solution due to Brouwer’s fixed point theorem. Due to the softmax, $F$ is neither a contraction nor is the resulting value iteration map monotonic unless $\beta_{h}<\beta_{h}^{1}$ for some $\beta_{h}^{1}>0$, so the solution can be non-unique in cyclic environments. Eqns. (4)–(9) also define a continuous self-map $G$ from a finite-dimensional bounded convex set $D$ of values $(Q_{r},\pi_{r},V^{e}_{h},X_{h},U_{r},V_{r})$ which is however not closed because $G$ is undefined for zero values in $Q_{r}$ or $X_{h}$. [^5] We could fix this by adding a small constant $\epsilon>0$ to $Q_{r}$ and $X_{h}$ before taking powers, which makes $G$ defined on a closed convex polytope as well, so that it must have a fixed point. Still, for $\beta_{r}=0$, there is a unique solution because $V^{e}_{h}$ is then the value function of a fixed policy, $U_{r}$ is independent of $V_{r}$, and so $V_{r}$ is also the value function of a fixed policy and reward function. Since $F,G$ are continuous in $\beta_{h},\beta_{r}$, we thus conjecture that homotopy / continuation methods can single out a unique “principal” solution for any $\beta_{h},\beta_{r}>0$ even in cyclic environments, as in Goeree, Holt, and Palfrey (2016).
@@ 3.1 Complex Multi-Agent Environments:
If $\mathcal{S}$ and $\mathcal{A}$ are large but $\mathcal{H}$ and $\mathcal{G}_{h}$ are not, the robot can use a tabular or approximate learning approach. [^6] If also $\mathcal{H}$ and $\mathcal{G}_{h}$ are large, $r$ could use single neural networks $Q^{m}_{\mathcal{H}},\pi_{\mathcal{H}},V^{m}_{\mathcal{H}},V^{e}_{\mathcal{H}},X_{\mathcal{H}}$ that take feature vectors describing $h$ and $g_{h}$ as additional inputs, and use samples $h,g_{h}$ to train these networks and to approximate the sums in eqns. (7) and (8).
@@
For each $h\in\mathcal{H},g_{h}\in\mathcal{G}_{h}$ separately, learn tables or neural network approximations of $Q^{m}_{h}$ and $\pi_{h}$. Generate samples $(s,g_{h},a,s^{\prime})$ using a slowly updated $\beta_{h}$-softmax policy $\pi_{h}(s,g_{h})$ based on $Q_{h}^{m}$ with a decreasing amount of additional exploration, the prior policy $a_{-h}\sim\mu_{-h}(s)$ for other humans, and, to eventually learn the minimum in (1), an $\epsilon$-greedy policy for $a_{r}$ based on the negative expected value $\operatorname*{\mathbb{E}}_{s^{\prime}\sim s,a}(U_{h}(s^{\prime},g_{h})+\gamma_{h}V^{m}_{h}(s^{\prime},g_{h}))$ with $\epsilon\to 0$. Then use expected SARSA targets on a time-scale faster than $\pi_{h}$.
@@
Based on the learned $\pi_{h}$, now aim to simultaneously learn tables or network approximations of $V^{e}_{h}$ and $X_{h}$ for all $h$, and of either $Q_{r}$ (DQN approach) or $V_{r}$ (actor-critic (AC) approach). Generate data samples $(s,g,a,s^{\prime})$ from rollouts using the fixed policies $\pi_{h}(s,g_{h})$, and either a $\beta_{r}^{\prime}$-softmax policy based on $Q_{r}$ with $\beta_{r}^{\prime}\nearrow\beta_{r}$, or a network approximation of $\pi_{r}$ trained on $V_{r}$ with entropy regularisation. Sample a new goal profile $g$ every $N_{g}$ steps. Instead of the direct expectation calculations of (4), (6), (7), then update the tables or networks via batched SGD using the update targets
@@ 4.1 Analysis of paradigmatic situations
We analyze the behavioral implications of our approach in several paradigmatic situations (details in the Supplement).
@@
If the robot can make binding commitments (e.g. by offering labelled buttons causing different behaviors), make secret plans to react to $h$’s actions in certain ways (e.g. by offering unlabelled buttons), or act without waiting for button presses, it will generally make commitments if it assumes $h$ to be sufficiently rational to use that information to act in ways that make $r$ do what $h$ actually wants. This is because $\pi_{h}$ does not depend on $\pi_{r}$ (encoding what $r$ plans) but on $\mathcal{A}_{r}$ (encoding what $r$ has committed to). Via such commitments, $r$ will establishing semantic connections between human actions (e.g., speech acts) and certain potentially complex behaviors of its own, to be able to act as an instruction-following assistant, somewhat similar to Reddy, Levine, and Dragan (2022).
@@
While an ‘empowerment’ maximizing robot would present humans with as many action options as possible, our robot will avoid overwhelming humans with too many options. If $r$ can choose to give $h$ any number $k$ of actions, each of which fulfils a different goal, it will choose $k\approx(e^{\beta_{h}}-1)/(\zeta-1)$ (noting that $\zeta>1$). This is because for larger $k$, $h$’s goal attainment probability will decrease due to bounded rationality so much that $h$’s effective power decreases despite the theoretical potential to fulfil more goals.
@@
The robot will sometimes ask for confirmation before obeying a command to perform an action because doing so decreases $h$’s error rate related to its bounded rationality. Also, if the action is irreversible, taking the action will remove $h$’s subsequent ability to revert that choice, so delaying action leaves $h$ with decision power for longer. However, $r$ will eventually obey because otherwise $h$ would not have that choice in the first place. As can be expected, the number of times $r$ asks back increases with larger $\gamma_{h}$ and $\gamma_{r}$, i.e., the more patient $h$ and $r$ are, and decreases with larger $\beta_{h}$. E.g., in a minimal model, with $\gamma_{h}=0.99$, $\gamma_{r}=0$, and an effective human error rate of 10%, $r$ will already ask back twice before obeying a command.
@@
The robot will tend to follow human social norms that generally foster goal achievement. This is because $r$ will model most $h$ as expecting most others to follow the norm ($\mu_{-h}$), will thus expect those $h$ to also follow the norm ($\pi_{h}$) since that increases $V^{m}_{h}(s,g_{h})$ for most $h$ and $g_{h}$. Thus $r$ will plan to follow the norm to prevent reducing those $h$’s power from harm or mis-coordination. If $r$ assumes humans to have internalized the norm into habits ($\pi^{0}_{h}$), it will even commit to following the norm to coordinate better with them, e.g. when passing each other in the street.
@@
If the robot can split a total amount $M$ of resources between $h_{1}$ and $h_{2}$, and for $h_{i}$ to have resources $m$ translates into having a power of $W_{h}=f(m)$ bits, then $r$ will generally prefer an equal split, at least if $f$ is linear [^7] E.g., a linear $f$ seems plausible if $M$ is money that can be spent for paying others to make independent choices in one’s favor. , concave, or not too convex, and will only prefer an unequal split or even full resource concentration if $f$ is very convex. Larger values of $\xi$ will make the split more equal.
@@
The robot might choose to make humans have incorrect beliefs $\mu_{-h}$ about each others’ behavior in situations where $r$ can do so in the first place and where correct beliefs would lower effective goal attainments $V^{e}_{h}$, e.g. when most strategic (quantal response) equilibria are bad and most “social optima” (in terms of total power $U_{r}$) are far from strategic equilibrium.
@@
If the robot cares for $h$’s future power ($\gamma_{r}\gg 0$) and can provide $h$ with the means to harm themselves, $r$ will trade off the temporary power increase from having these means against the possible later disempowerment from harm. So $r$ will provide the means only if it believes that $h$ is sufficiently rational and that $h$ will most probably not actually harm themselves.
@@
If the robot has a pause and a destroy button and might disable either, it will generally enable only the pause button to prevent partially disempowering $h$ permanently by not being able to assist $h$ when destroyed. Only if the robot thinks it is very unlikely that $h$ will use the destroy button will it enable that button to give $h$ this additional power. If $r$ considers itself very empowering, it might even disable the pause button.